A sample neon has a volume of 5400ml at 29.5 C. At what temperature would the gas occupy a volume of 1000ml?


1
Expert's answer
2021-04-26T03:01:19-0400

V1/T1 = V2/T2

5400/(29.5 + 273) = 1000/T2

5400/302.5 = 1000/T2

17.85T2 = 1000

= 56K

= -217°C